OpenStudy (anonymous):

I need to find the equation of a parabola that has a vertex of (-3,0) and is opening upwards.

OpenStudy (loser66):

I try to help, but forgot the formula of parabola, do you remember?

OpenStudy (anonymous):

I have not learned that actually all I know is the factoring method and quadratic formula

OpenStudy (loser66):

ok, google now, both us , OK? and then we compete, play or not?

OpenStudy (anonymous):

I tried googling how to do this but no answer

OpenStudy (loser66):

not how to do, we google " what is the vertex formula of a parabola" .

OpenStudy (anonymous):

Ok

OpenStudy (loser66):

3,2,1 get go

OpenStudy (loser66):

hey, I got it first, I won, so, you have to find out the meaning of the formula. ok?

OpenStudy (anonymous):

I have it but I am confused on how to use it

OpenStudy (loser66):

hehe, loser, now your turn, explain me, what h,k mean?

OpenStudy (anonymous):

Ummm. This website has nothing on how to go from a vertex point to an equation

OpenStudy (loser66):

so, what you have from your website?

OpenStudy (anonymous):

I had the exact same website but it is not what I am asking

OpenStudy (anonymous):

How do I go from a vertex (-3,0) To an entire equation of the parabola ?

OpenStudy (loser66):

hhmmm I don't know either, ok, get help from other, you tag someone.

OpenStudy (loser66):

@e.mccormick

OpenStudy (loser66):

@Mertsj

OpenStudy (anonymous):

Does anyone know

OpenStudy (campbell_st):

the general form is \[y =a(x -h)^2 + k\] where (h, k) is the vertex and a is a constant so you have \[y = a( x - 3)^2 + 0\]

OpenStudy (campbell_st):

in your question I'd assume a = 1

OpenStudy (anonymous):

I do not know what a is

OpenStudy (loser66):

hold on!! it's (x+3) not (x-3)

OpenStudy (loser66):

since our h is -3 , am I right?

OpenStudy (anonymous):

idk what the a is

OpenStudy (anonymous):

H should be -3?

OpenStudy (campbell_st):

well then its \[y = a( x + 3)^2 + 0\] vertex ( -3 , 0) and let a = 1

OpenStudy (loser66):

yes, !! in general , whatever a is, we must have a positive to have the parabola upward.

OpenStudy (anonymous):

so now what?

OpenStudy (campbell_st):

thats partly correct as \[y = ( x+ 3)^2 + 0....and.....y = 4(x + 3)^2 + 0\] have the same vertex but are different graphs...

OpenStudy (campbell_st):

well all you need to do is simplify \[y = (x + 3)^2 \] just distribute

OpenStudy (anonymous):

\[x ^{2} + 9?\]

OpenStudy (campbell_st):

you are missing the middle term 6x and its y = ......

OpenStudy (anonymous):

Wait!!

OpenStudy (loser66):

I hand off, @Plungedcashew good luck, you have a very high smartscore person helps you. the game is not fun anymore

OpenStudy (anonymous):

I get to match parabolas to equations

OpenStudy (anonymous):

And one of the possible answers is actually y=(x+3)2

OpenStudy (campbell_st):

no more work to do.

OpenStudy (anonymous):

Wait so..

OpenStudy (anonymous):

Why is it x+3?

OpenStudy (anonymous):

Oh I got it!

OpenStudy (anonymous):

Thanks sooo much man. How do I give you a triple AAA rating?

OpenStudy (campbell_st):

ok.... here is the important bit the vertex form of a parabola is \[y = a(x - h)^2 + k\] (h, k) is the vertex. In your question the vertex is (-3, 0) so h = -3 and k = 0 substitute them for the answer.

OpenStudy (anonymous):

Thanks man. i am glad you came along for me!